OpenStudy (anonymous):

Two spheres of the same density have a ratio of 4 to 9 in surface area. If the small sphere weighs 10kg, what does the other sphere weigh?

OpenStudy (ganpat):

Surface Area = 4 * Pi * r2 A1 / A2 = r1 sq / r2 sq 4 / 9 = r1 sq / r2 sq taking sq rt... So r1 / r2 = 2/3 Volume of sphere = 4/3 * pi * r3 So for two spheres... V1 / V2 = r1 cube / r2 cube = (2 / 3 ) power of cube = 8/ 27 Now, Density is same for both spheres, and density = Mass / Volume.. M1 / V1 = M2 / V2 Given, M1 = 10 Kg so M2 = (10 * 27) / 8 = 33.75 Kg... is that right answer ???

OpenStudy (anonymous):

I'll find out tomorrow :S Hahaha Thanks

OpenStudy (ganpat):

Method is definitely right, might be calculation error.. and , do not blame me, if something goes wrong.... :D All the Best !!!!!